1.     Consider the following model of a hypothetical market for MSW management services:

 

                                   MPB   =         25 − 2Q           

                                   MPC    =          4 + Q

                                   MEC    =         0.5Q,

           where Q is the number of trash containers per household per month.

A.     Quantitatively determine the effect of the resource misallocation due to: (12 marks)

a.     the presence of the negative externality, and

b.     the use of a flat fee pricing system in the presence of a negative externality.

B.     Determine the dollar value of a waste-end charge that would restore efficiency to this market. Explain your answer intuitively. (3 marks)



1
Expert's answer
2021-12-12T16:39:07-0500

A

i)

Under equilibrium, demand is equal to supply,

25 − 2Q =  4 + Q

Therefore, Q = 7


At price p the equilibrium quantity is 7

price p = 25 - 2(7)

= 11

The price is 11 when externality is not incorporated in the supply function.

When cost of externality is considered,

S' = MPC +MEC

S' = 4 + Q + 0.5Q

= 4 + 1.5Q

Therefore, demand = S'

25 - 2Q = 4 + 1.5Q

Q' = 6

When externality is considered, the output reduces to 6

Resource misallocation of 1 will be provided to the market when the negative externality is not considered.

Flat fee is a fee that is considered despite the quantity.

The cost of additional waste management is zero.

MPC = 0

Under Equilibrium MPB = MPC

25 -2Q = 0

Q = 12.5

Therefore, resource misallocation is 12.5 - 7 = 5.5

B

Waste end charge is the cost that is charged when the waste is disposed and it is depends on the amount of waste that is generated. To restore efficiency, dollar value of waste minus end charge should be equal to MPC -MEC at a socially optimum level of output. to cover both the private cost of waste plus the cost of externality. Socially optimum level of output is 6. Marginal social cost at this point is 13. Therefore, the fee should be 13 per unit.